OpenStudy (anonymous):

How many liters each of a 45% acid solution and a 75% acid solution must be used to produce 60 liters of a 70% acid solution? (round to two decimal places if necessary)

OpenStudy (anonymous):

x+y=60 .45x+.75y=60(.70)

OpenStudy (amistre64):

a + b = t a = t-b a(x) + b(y) = t(z) (t-b)(x) + b(y) = t(z) tx -bx +by = tz b(y-x) = t(z-x) b = t(z-x) ----- (y-x) b = 60(.70-.45) at 45% ---------- (.45-.75)

OpenStudy (amistre64):

got my numbers backwards on the bottom :) b = 60(.70-.45) at 45% ---------- (.75-.45)

OpenStudy (amistre64):

50 at 75% ... seems better; still new at it :)

OpenStudy (anonymous):

and then 10 at 75%?

OpenStudy (amistre64):

yes; but (50) at 75% and (10) at 45%